OpenStudy (anonymous):

Which equation below is not an equation that represents the graph shown below?

OpenStudy (anonymous):

CHOICES: A. y - 3 = 2(x - 6) B. y + 3 = 1/2 (x + 6) C. 2y = x D. x - 2y = 0

OpenStudy (anonymous):

its a

OpenStudy (ash2326):

The line passes through the point (0, 0) and the point (2, 1) Check @ashleyxo1 Which one satisfies these two conditions?

OpenStudy (anonymous):

D

OpenStudy (ash2326):

Correct :D

OpenStudy (anonymous):

Thanks! (: